# JSJ splittings for all Artin groups

### v1

@@ Abstract.
We prove that an Artin group splits over infinite cyclic subgroups if and only if its defining graph has a separating vertex, and explicitly construct a JSJ decomposition over infinite cyclic subgroups for all Artin groups. This is then used to show that, if two Artin groups are isomorphic, then they have the same set of big chunks, that is, maximal subgraphs without separating vertices. We also deduce acylindrical hyperbolicity for the automorphism groups of many Artin groups, partially answering a question of Genevois in the case of Artin groups.
@@
Key words. Artin groups, Bass-Serre theory, JSJ decompositions, isomorphism problem.
@@ Introduction
In this paper we investigate when an Artin group splits over $\mathbb{Z}$, meaning that it admits a graph of group decomposition with infinite cyclic edge groups. Such splittings assumed a central role in the solution of the isomorphism problem for hyperbolic groups (see [Sel95, Bow98, RS97, DG11] among many others); in particular, “maximal” splittings called JSJ decompositions were the key tool for understanding automorphisms of hyperbolic groups. In a similar fashion, we also obtain results related to the isomorphism problem for Artin groups, and establish some geometric properties of automorphism groups of Artin groups (see Theorem B and Corollary E, respectively).
@@
One of the most widely open questions for Artin groups is the isomorphism problem, which asks for an algorithm which takes two labelled graphs $\Gamma$ and $\Gamma^{\prime}$ as input, and determines if they yield isomorphic Artin groups. This question, which specialises Dehn’s Isomorphism Problem to the family of Artin groups, is solved when one restricts to certain subclasses, including RAAGs [Dro87], spherical-type [Par04], and large-type [Vas22], but there is no unified approach. However, isomorphic Artin groups must admit the same splittings over $\mathbb{Z}$, as this is an algebraic property. In this spirit, if $A_{\Gamma}\cong A_{\Gamma^{\prime}}$, then one can hope to use Theorem A to recover some decompositions of $\Gamma$ and $\Gamma^{\prime}$ into maximal connected subgraphs without separating vertices, which we call big chunks. We do just that in Theorem 5.6, of which we give a streamlined version here. For a labelled graph $\Gamma$, let $\mathcal{BC}(\Gamma)$ be the collection of its big chunks.
@@
JSJ decompositions, trees of cylinders, and deformation spaces
@@
Our JSJ decomposition is not unique, as different Artin generating sets of the same Artin group may yield non-isomorphic JSJ decompositions. There is a procedure, due to Guirardel and Levitt, to obtain a canonical tree, called the tree of cylinders, from the JSJ tree. This tree is left invariant by the automorphism group, and the action turns out to be acylindrical, allowing us to affirmatively answer [Gen19, Question 1.1] for any Artin group whose defining graph contains a separating vertex.
Corollary E (see Corollary 4.6).
@@
We remark that the assumption that $A_{\Gamma}$ is torsion-free is conjecturally trivial, since all Artin groups are expected to be torsion-free. By work of Fournier-Facio and Wade, Corollary E implies that an Artin group $A_{\Gamma}$ as in the statement has an infinite-dimensional space of $\operatorname{Aut}(A_{\Gamma})$-invariant quasimorphism (see [FFW23, Theorem E]). In turn, this has applications to construct unbounded $\operatorname{Aut}(A_{\Gamma})$-invariant word norms on $A_{\Gamma}$ (see [FFW23, Corollary C]).
@@
In Section 3 we produce a JSJ decomposition, thus proving Theorem C (see Definition 3.7 for the construction of the splitting, and Theorem 3.9 for the proof that it yields a JSJ decomposition). Then in Section 4 we combine the JSJ with the machinery of [GL11] to prove Corollary E on acylindrical hyperbolicity of automorphism groups of certain Artin groups (see Corollary 4.6).
@@ Acknowledgements
We thank Yassine Guerch and Gilbert Levitt for suggesting the application to acylindrical hyperbolicity of automorphism groups; we are also grateful to Kaitlin Ragosta and Ruta Sliazkaite for insightful discussions. A special thanks goes to our supervisors Laura Ciobanu, Alessandro Sisto, and Alexandre Martin for comments on a first draft of this document.
@@ 4.2. Trees of cylinders
Let $G$ be a group whose torsion subgroups have uniformly bounded cardinality. Let $(T,\Omega)$ be a JSJ-tree over virtually cyclic subgroups of $G$ which is not a point. Then $G$ acts minimally and acylindrically on $T_{c}$, and the action extends to an isometric action of $\operatorname{Aut}(G)$.
@@
We already noticed that the $G$-action is minimal, and it extends to $\operatorname{Aut}(G)$ by [GL11, Corollary 4.10], so we are left to prove acylindricity. Let $p,q\in V(T_{c})$ be at distance at least $6$, and we claim that $\operatorname{Stab}_{\Omega_{c}}\left(p\right)\cap\operatorname{Stab}_{\Omega_{c}}\left(q\right)$ has uniformly bounded cardinality. Since the tree is bipartite, we can find $p^{\prime},q^{\prime}\in V_{0}$ which lie on a geodesic $[p,q]$ and are at distance at least $4$. Since any element fixing $p$ and $q$ must also fix $p$ and $q^{\prime}$, it is enough to bound the cardinality of $\operatorname{Stab}_{\Omega_{c}}\left(p^{\prime}\right)\cap\operatorname{Stab}_{\Omega_{c}}\left(q^{\prime}\right)$. Since $p^{\prime},q^{\prime}\in V_{0}$, the latter equals $H\coloneq\operatorname{Stab}_{\Omega}\left(p^{\prime}\right)\cap\operatorname{Stab}_{\Omega}\left(q^{\prime}\right)$, which is therefore virtually cyclic as it is contained inside the stabiliser of some edge of $T$. If $H$ is finite then we are done, as torsion subgroups of $G$ have uniformly bounded cardinality. Otherwise $H$ is commensurable to all edge stabilisers on a $T$-geodesic between $p^{\prime}$ and $q^{\prime}$, which therefore belong to the same cylinder, violating the fact that $\operatorname{d}_{T_{c}}(p^{\prime},q^{\prime})\geq 4$. ∎
@@ 4.3. Application to automorphism groups of Artin groups
Now let $(T,\Omega)$ be the Bass-Serre tree of the JSJ splitting $J(\Gamma)$ from Definition 3.7, which is not a point since $\Gamma$ is not a single big chunk. By Proposition 4.5, $A_{\Gamma}$ has a minimal, acylindrical action on the tree of cylinders $T_{c}$ which extends to $\operatorname{Aut}(A_{\Gamma})$. We take a small detour to prove the following:
@@
Finally, $\operatorname{Aut}(A_{\Gamma})$ is acylindrically hyperbolic since the $\operatorname{Aut}(A_{\Gamma})$-action on $T_{c}$ satisfies the requirements of [GH21, Proposition 3.8]:
@@ 5. Isomorphism invariance of big chunks
    If $\Lambda$ is an even leaf, then so is $\varphi_{\#}(\Lambda)$.

### v2

@@ Abstract.
We prove that an Artin group splits over infinite cyclic subgroups if and only if its defining graph has a separating vertex, and explicitly construct a JSJ decomposition over infinite cyclic subgroups for all Artin groups. This is then used to show that, if two Artin groups are isomorphic, then they have the same set of big chunks, that is, maximal subgraphs without separating vertices. We also deduce acylindrical hyperbolicity for the automorphism groups of many Artin groups, partially answering a question of Genevois in the case of Artin groups. As a consequence, this produces new families of Artin groups with the $R_{\infty}$ property.
@@
Key words. Artin groups, Bass-Serre theory, JSJ decompositions, isomorphism problem, acylindrical hyperbolicity.
@@ Introduction
In this paper we investigate when an Artin group splits over $\mathbb{Z}$, meaning that it admits a graph of group decomposition with infinite cyclic edge groups. Such splittings assumed a central role in the solution of the isomorphism problem for hyperbolic groups (see [Sel95, Bow98, RS97, DG11] among many others); in particular, “maximal” splittings called JSJ decompositions were the key tool for understanding automorphisms of hyperbolic groups. In a similar fashion, we also obtain results related to the isomorphism problem for Artin groups, and establish some geometric properties of automorphism groups of Artin groups (see Theorem B and Theorem E, respectively).
@@
One of the most widely open questions for Artin groups is the isomorphism problem, which asks for an algorithm which takes two labelled graphs $\Gamma$ and $\Gamma^{\prime}$ as input, and determines if they yield isomorphic Artin groups. This question, which specialises Dehn’s Isomorphism Problem to the family of Artin groups, is solved when one restricts to certain subclasses, including RAAGs [Dro87, Bau81], spherical-type [Par04], and large-type [Vas22], but there is no unified approach. However, isomorphic Artin groups must admit the same splittings over $\mathbb{Z}$, as this is an algebraic property. In this spirit, if $A_{\Gamma}\cong A_{\Gamma^{\prime}}$, then one can hope to use Theorem A to recover some decompositions of $\Gamma$ and $\Gamma^{\prime}$ into maximal connected subgraphs without separating vertices, which we call big chunks. We do just that in Theorem 5.6, of which we give a streamlined version here. For a labelled graph $\Gamma$, let $\mathcal{BC}(\Gamma)$ be the collection of its big chunks.
@@
JSJ decompositions for Artin groups
@@
An example of the JSJ decomposition for an Artin group. There is one white vertex for every separating vertex, and one black vertex for every big chunk. Even leaves further split as HNN extensions (if the label is two), or amalgamated products otherwise.
@@
Acylindrical hyperbolicity of automorphism groups and consequences
Our JSJ decomposition is not unique, as different Artin generating sets of the same Artin group may yield non-isomorphic JSJ decompositions. By work of Guirardel and Levitt [GL11], there is a procedure that takes as input the JSJ tree and produces a canonical, $\operatorname{Aut}(A_{\Gamma})$-invariant tree, called the tree of cylinders. Under suitable conditions on $\Gamma$, we can use the tree of cylinders to prove that $\operatorname{Aut}(A_{\Gamma})$ is acylindrically hyperbolic:
Theorem E (see Corollary 4.6).
@@
We remark that the assumption that $A_{\Gamma}$ is torsion-free is conjecturally trivial, since all Artin groups are expected to be torsion-free. The above Corollary affirmatively answers [Gen19, Question 1.1] for any Artin group whose defining graph contains a separating vertex.
An Artin group as in Theorem E is centerless, as it is acylindrically hyperbolic and torsion-free (see [Osi16, Corollary 7.2]). A centerless group with acylindrically hyperbolic automorphism group has several remarkable features, including the so-called $R_{\infty}$ property:
Corollary F (of [IMSW25, Corollary 8.1.4]).
An Artin group $A_{\Gamma}$ as in Theorem E has property $R_{\infty}$.
Recall that, given a group $G$ and an automorphism $\varphi\in\operatorname{Aut}(G)$, two elements $g,h\in G$ are in the same $\varphi$-twisted conjugacy class if there exists $k\in G$ such that $\varphi(k)gk^{-1}=h$. Furthermore, $G$ has property $R_{\infty}$ if every automorphism $\varphi\in\operatorname{Aut}(G)$ admits infinitely many twisted conjugacy classes. The $R_{\infty}$ property finds its origin in Nielsen fixed point theory, where under certain assumptions it relates to the number of fixed points of homeomorphisms of a space with fundamental group $G$. Among Artin groups, the $R_{\infty}$ property had been previously established for some spherical-type and affine-type Artin groups [FG10, DLGO21, CS22], for several classes of large-type Artin groups [Juh12, SV24], and for RAAGs [DS21, Wit23].
We also list some other consequences of Theorem E, which exhibit the existence of several non-trivial $\operatorname{Aut}(A_{\Gamma})$-invariant objects:
Corollary G.
An Artin group $A_{\Gamma}$ as in Theorem E admits:
  • an infinite simple characteristic quotient [CFF23, Corollary 1.4];
  • an infinite-dimensional space of $\operatorname{Aut}(A_{\Gamma})$-invariant quasimorphism [FFW23, Theorem E];
  • unbounded $\operatorname{Aut}(A_{\Gamma})$-invariant word norms [FFW23, Corollary C].
The JSJ deformation space
@@
In Section 3 we produce a JSJ decomposition, thus proving Theorem C (see Definition 3.7 for the construction of the splitting, and Theorem 3.9 for the proof that it yields a JSJ decomposition). Then in Section 4 we combine the JSJ with the machinery of [GL11] to prove Theorem E on acylindrical hyperbolicity of automorphism groups of certain Artin groups (see Corollary 4.6).
@@ Acknowledgements
We thank Yassine Guerch and Gilbert Levitt for suggesting the application to acylindrical hyperbolicity of automorphism groups, and Francesco Fournier-Facio for pointing us towards its corollaries. We are also grateful to Kaitlin Ragosta and Ruta Sliazkaite for insightful discussions. A special thanks goes to our PhD supervisors Laura Ciobanu, Alessandro Sisto, and Alexandre Martin for comments on a first draft of this document.
@@ 4.2. Trees of cylinders
Let $G$ be a group whose torsion subgroups have uniformly bounded cardinality. Let $(T,\Omega)$ be a JSJ-tree over virtually cyclic subgroups of $G$ which is not a point. Then:
  • $G$ acts minimally and acylindrically on $T_{c}$;
  • The centre $Z(G)$ acts trivially;
  • The $G/Z(G)$-action extends to $\operatorname{Aut}(G)$.
@@
We already noticed that the $G$-action is minimal, so we now prove acylindricity. Let $p,q\in V(T_{c})$ be at distance at least $6$, and we claim that $\operatorname{Stab}_{\Omega_{c}}\left(p\right)\cap\operatorname{Stab}_{\Omega_{c}}\left(q\right)$ has uniformly bounded cardinality. Since the tree is bipartite, we can find $p^{\prime},q^{\prime}\in V_{0}$ which lie on a geodesic $[p,q]$ and are at distance at least $4$. Since any element fixing $p$ and $q$ must also fix $p$ and $q^{\prime}$, it is enough to bound the cardinality of $\operatorname{Stab}_{\Omega_{c}}\left(p^{\prime}\right)\cap\operatorname{Stab}_{\Omega_{c}}\left(q^{\prime}\right)$. Since $p^{\prime},q^{\prime}\in V_{0}$, the latter equals $H\coloneq\operatorname{Stab}_{\Omega}\left(p^{\prime}\right)\cap\operatorname{Stab}_{\Omega}\left(q^{\prime}\right)$, which is therefore virtually cyclic as it is contained inside the stabiliser of some edge of $T$. If $H$ is finite then we are done, as torsion subgroups of $G$ have uniformly bounded cardinality. Otherwise $H$ is commensurable to all edge stabilisers on a $T$-geodesic between $p^{\prime}$ and $q^{\prime}$, which therefore belong to the same cylinder, violating the fact that $\operatorname{d}_{T_{c}}(p^{\prime},q^{\prime})\geq 4$.
Moving to the second bullet, every element in the centre fixes every cylinder, as it maps edges of $T$ to edges with the same stabilisers. As every vertex in $V_{0}$ belongs to at least two cylinders, this means that the centre acts trivially on $T_{c}$. Then the $G/Z(G)$-action, which we identify with $\operatorname{Inn}(G)\leq\operatorname{Aut}(G)$, extends to $\operatorname{Aut}(G)$ by [GL11, Corollary 4.10]. ∎
@@ 4.3. Application to automorphism groups of Artin groups
Now let $(T,\Omega)$ be the Bass-Serre tree of the JSJ splitting $J(\Gamma)$ from Definition 3.7, which is not a point since $\Gamma$ is not a single big chunk. By Proposition 4.5, $A_{\Gamma}$ has a minimal, acylindrical action on the tree of cylinders $T_{c}$, which extends to $\operatorname{Aut}(A_{\Gamma})$. We take a small detour to prove the following:
@@
Finally, to prove that $\operatorname{Aut}(A_{\Gamma})$ is acylindrically hyperbolic it is enough to verify the requirements of [GH21, Proposition 3.8]:
@@ 5. Isomorphism invariance of big chunks
The new splitting has one white vertex for every separating vertex and one black vertex for every big chunk. The vertex group of a black vertex is now the full big chunk parabolic, also for even leaves.